OpenStudy (anonymous):

What is the value of g[f(4)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5?

OpenStudy (anonymous):

f(4) = 9 g(9) = 13

OpenStudy (diyadiya):

f(x) = 2x + 1 plug in 4 in place of x f(4)=2(4)+1=9 g( f(4) )=g(9) g(x) = 2x - 5 x=9 here g(9)=2(9)-5=18-5=13

OpenStudy (diyadiya):

Understood?

OpenStudy (anonymous):

yes

OpenStudy (diyadiya):

=)